OpenStudy (lina17):

help please solve for x: -5|x+1|=10

alones (alones):

you have any answer choice to your question??

OpenStudy (lina17):

A. x=0 b. x=-3 and x=1 c. x=-1 and x=3 d. no solution

alones (alones):

Oh okay, let's solve your equation step-by-step. −5(|x−1|)=10 Step 1: Divide both sides by -5. −5(|x−1|)−5=10−5 |x−1|=−2 Step 2: Solve Absolute Value. |x−1|=−2 No solutions. (Absolute value cannot be less than 0.) Answer: No solutions. Hope that helps :)

OpenStudy (lina17):

ok thanks i got no solution for my answer too

alones (alones):

Great Job<33 :)

OpenStudy (lina17):

can you help me on one more

alones (alones):

umm i think i can :)

OpenStudy (lina17):

\[solve for x: |2x+6|-4=20\]

OpenStudy (lina17):

i got no solution em i right

alones (alones):

wait you got no solution hmmm how??

alones (alones):

I got x=9 or x=−15

OpenStudy (lina17):

omg im so stupid lol i did the steps wrong

OpenStudy (lina17):

thanks

alones (alones):

no your not stupid I mean I hopue Your not hehe ;]

OpenStudy (lina17):

lol im not

alones (alones):

thanks :)